# Strange stars admixed with dark matter: equiparticle model in a two fluid approach

### v2

@@ Abstract
In this work we explore the possible scenario of strange stars admixed with fermionic or bosonic dark matter. For the description of the “visible” sector, we use a specific phenomenological quark model that takes into account in-medium effects for the quark masses through a suitable baryonic density dependence, in which the free parameters are chosen from the analysis of the stellar matter stability window (parametrizations presenting lower energy per baryon than iron nuclei). In the case of the dark sector, we investigate the predictions of fermionic and bosonic models. In the former we consider a spin $1/2$ dark particle, and the latter is described by a dark scalar meson. Both models present a repulsive vector interaction, particularly important for the bosonic model since it helps avoiding the collapse of the star due to the lack of the degeneracy pressure. Our results point out to possible stable strange stars configurations in agreement with data from PSR J0030+0451, PSR J0740+6620, NICER and HESS J1731-347. As another feature, we find stars with dark matter halo configurations for lower values of the dark particle mass.
@@ I Introduction
Stars are some of the most studied astrophysical systems in the universe. The detailed composition of these compact objects, as well as their structure and dynamics, are still not completely understood by astrophysicists and, for decades, it has been subject of intense research encompassing many areas of physics such as thermodynamics, quantum field theory, general relativity, and nuclear reactions, for instance. Features of the strongly interacting matter can also be investigated due to the extreme conditions present in the core of the stars, region expected to be at very high density regime. The input used to correctly describe the compressed stellar matter (CSM) comes from models where hadrons (baryons and mesons) are the proper degrees of freedom. Nevertheless, there is also another branch of study that considers quarks as the building blocks in models from which equations of state are derived and applied to the CSM formulation. This approach is based on the quark stars hypothesis proposed long time ago in [1, 2], and later on supported by Bodmer [3] and Witten [4] who established the so called Bodmer-Witten conjecture: quark matter consisting of $u$, $d$, and $s$ quarks in beta equilibrium can present lower energy per baryon than hadronic matter. Evidences for quark matter composition in cores of massive stars were recently discussed in Refs. [5, 6].
@@
Both hadronic and phenomenological quark models, constructed for the aforementioned purpose, can be tested against a large number of new astrophysical observational measurements, recently emerged due to (i) the detection of gravitational waves, coming from a binary system with its respective electromagnetic counterparts [7], made by the LIGO and Virgo Collaboration [8, 9, 10], (ii) the data concerning the massive millisecond pulsars PSR J0030+0451 [11, 12] and PSR J0740+6620 [13, 14], provided by the NASA’s Neutron star Interior Composition Explorer (NICER) x-ray telescope based on the International Space Station [15], and (iii) the first simultaneous measurements of mass, radius, and surface temperature of the HESS J1731-347 compact star [16].
@@
Recently many researchers have been dedicated efforts to better comprehend the physics of these objects, that may present another mysterious component as well, named as dark matter (DM) [17, 18, 19]. The main evidences for the existence of non-luminous matter come from the analysis of galaxies rotation curves, and from the use of gravitational lensing methods [20, 21], that lead to visible galaxy clusters masses of around 10% to 20%, being the remaining mass due to DM. Measurements of the anisotropy of cosmic microwave background also suggest presence of DM [22, 23, 24, 25, 26, 27]. It is a consensus that the universe, as we know nowadays, is 27% composed by DM, 68% of dark energy, and 5% of ordinary matter.
@@
In this work we study the possibility of existence of strange stars admixed with fermionic and bosonic dark matter, in agreement with the astrophysical constraints mentioned before, by using a particular effective quark model in the “visible” sector, the equiparticle (EQP) model in our case, in which constituent quark masses ($m_{u,d,s}$) are medium dependent, more specifically density dependent in the case of stellar matter at $T=0$. Temperature effects on $m_{u,d,s}$ can be found in [28, 29, 30, 31], and in [32], modified density dependencies of these quantities are proposed. The thermodynamical inconsistency of the original model was solved in [33] through a suitable connection between the quark Fermi momentum and the respective effective chemical potential. In [34], some of us have improved this model in order to take into account deconfinement effects through the introduction of the traced Polyakov loop as the order parameter of the confinement/deconfinement first order phase transition. Such a formulation was called PEQM model.
@@
Here we adopt the perspective of the two fluid approach, that only considers gravitational interaction between DM and quark matter. The calculations and results coming from this investigation, all of them in units of $\hbar=c=G=1$, are organized in this paper as follows. In Sec. II we present the main quantities related to effective EQP model, and establish the conditions for the physical stability window. In Sec. III we detail the main equations of sate used for the dark sector of our analysis, describing the fermionic and bosonic DM models. Results related to strange stars obtained from this formalism are shown to the reader in Sec. IV. Finally, our final remarks are given in Sec. V.
@@ II.1 Formalism
The EQP model was proposed in [33] as a baryonic density ($\rho_{b}$) and temperature ($T$) dependent model. However, here we focus on the zero temperature regime, as done in [35] and [34], for the study of cold quark stars. For this reason, the formalism of the model will be presented for this case. For a infinitely large system composed by up ($u$), down ($d$) and strange ($s$) quarks, the equations of state are obtained from fundamental thermodynamical relations. For the pressure we have
@@
where $m_{I}$ is the density dependent quark mass term. The energy density reads
@@
Following up on this, for the EQP model the quark masses is constructed as being density dependent as follows:
@@
where $m_{i0}$ is the current mass of the quark $i$. The interaction effects between the quarks are contained in the parameters $C$ and $D$, which are chosen from a stability window which will be described later. When working with compact objects, there is an interest of achieving higher stellar masses in order to satisfy the observational constraints. For this reason, the parameter $C$ plays an import role highlighting the necessity of incorporating this parameter into the mass scaling derived from [33] and presented in Eq. (10).
@@ II.2 Stellar matter
Given the interest in studying compact objects it is necessary to establish the conditions for the matter present in its interior. At zero temperature and high density regime, stellar matter in pure quark or hybrid stars [41, 42, 43, 44, 45, 46, 47, 48] can be described by effective quark models.
The most powerful mechanism of neutrino emission in stars is the Urca process [49, 50], responsible for their cooling since in this process stars lose a significant amount of thermal energy. The rate of neutrino emission influences how quickly the star cools down from its initial high temperature. After neutrinos are emitted from the star, its internal conditions undergoes through some changes, where the beta equilibrium takes place. This condition is given by $\mu_{u}+\mu_{e}=\mu_{d}=\mu_{s}$ in quark stars, which are the chemical potentials of the particles involved. In therms of the previously discussed model, this relation can be written as $\mu_{u}^{*}+\mu_{e}=\mu_{d}^{*}=\mu_{s}^{*}$, where the chemical potentials of the quarks are now the effective ones. The electron ($e$) does not have an effective chemical potential because its mass is constant. This equilibrium condition ensures that the star remains stable and maintains its composition over cosmic timescales. Furthermore, another condition has to be satisfied in compact stars, namely, the charge neutrality condition, given here by $\frac{2}{3}\rho_{u}-\frac{1}{3}\rho_{d}-\frac{1}{3}\rho_{s}-\rho_{e}=0$. These conditions exist due to the presence of leptons (electrons in this case), where weak interactions happen such as $d,s\leftrightarrow u+e+\bar{\nu}_{e}$. Because of the electrons presence, stellar matter is described by the sum of quarks and electrons contribution. Considering the (massless) electrons contribution, the total energy density and pressure of the part of the system composed by “visible” matter are
@@ II.3 Stability window
For the discussion of strange quark matter (SQM) in the context of quark stars, one has to consider the Bodmer-Witten hypothesis [3, 4]. It suggests that under certain conditions, such as at high densities or pressures, SQM may be the true ground state of matter rather than normal nuclear matter. Therefore, in order to establish the stability of SQM we need to investigate it against nuclear matter, which can be done through evaluating the minimum of the energy per baryon, $(E/A)_{\mbox{\tiny min}}$. This evaluation is then classified as follows: (i) SQM is stable when the minimum of the energy per baryon is lower than the binding energy of 56Fe (because it is the heaviest stable nucleus found in nature), i.e., $(E/A)_{\mbox{\tiny min}}\leqslant 930$ MeV; (ii) SQM is considered to be meta-stable when the minimum of the energy per baryon is higher than 930 MeV and lower than 939 MeV, which is the nucleon mass, i.e., 930 $930\mbox{ MeV}<(E/A)_{\mbox{\tiny min}}\leqslant 939$ MeV; (iii) SQM is unstable only if the minimum of the energy per baryon is higher than 939 MeV, i.e, $(E/A)_{\mbox{\tiny min}}>939$ MeV. In the case of stellar matter, we have $E/A=\mathcal{E}_{\mbox{\tiny vis}}/\rho_{b}$.
@@
In the figure we observe different regions of stability for SQM. The pink one corresponds to the absolutely stable state of SQM, which is our area of interest. Below this area we find the green region where two-flavor quark matter (2QM) is stable and where the SQM is forbidden to happen. In the orange region the meta-stable state of SQM is expected to occur, and in blue we have the unstable SQM. It is important to mention that this analysis takes in consideration the data provided by PDG [51] regarding the ranges for the current quark masses, $m_{u0}=2.16^{+0.49}_{-0.26}$, $m_{d0}=4.67^{+0.48}_{-0.17}$ MeV and $m_{s0}=93.4^{+8.6}_{-3.4}$ MeV. For stellar matter the chosen values are $m_{u0}=2.16$ MeV , $m_{d0}=5.15$ MeV, and $m_{s0}=90$ MeV. For the pair $(C,D)$ we take $C=0.81$ and $D^{1/2}=127$ MeV, since it produces more massive stars when only quark matter is considered.
@@ II.4 Polyakov equiparticle model
The Polyakov equiparticle (PEQP) model was proposed in [52, 34] to study the confinement/deconfinement phase transition, possible to happen inside strange stars. In that article, some of us perform a few modifications in the original EQP model, which will be briefly described in this section. In this improved model, gluons and its corresponding strong interaction dynamics is phenomenologically described by the traced Polyakov loop, $\Phi$. When $\Phi=0$, the quarks are confined and with $\Phi\rightarrow 1$ they achieve the deconfined phase. In the NJL model [53, 54, 55, 56, 57, 58], this quantity was introduced, at finite temperature [59], to generate the so-called Polyakov-Nambu-Jona-Lasinio (PNJL) model [60, 61, 62, 63, 64, 65, 66]. At first, the PNJL model loses its characteristics of describing the confinement/deconfinement phase transition at $T=0$, but in [67] the authors were able to solve this problem by taking the $b_{2}(T)$ function, present in the Polyakov potential, and modifying it to become a function that is also dependent on the quark density. Later on, the PNJL model suffered another improvement done in [68, 69, 70], where the motivation was to ensure that the coupling constants would vanish at the deconfinement phase, i.e. at $\Phi=1$.
All these observations concerning the improvement of the original NJL model were also implemented in the EQP model. More specifically, the parameters $C$ and $D$ were modified as follows, $C\rightarrow C^{\prime}(C,\Phi)=C(1-\Phi^{2})$, and $D\rightarrow D^{\prime}(D,\Phi)=D(1-\Phi^{2})$, which lead us to the a new equation for the quark masses,
@@ III Dark matter models
With regard to the dark sector of our approach, we choose to investigate the capability of the DM model in satisfying the recent astrophysical constraints when admixed to the previously described quark matter, by implementing a fermionic and a bosonic model. Concerning the first one, we adopt, as starting point, a Lagrangian density presenting a kinetic term (Dirac Lagrangian density) of a single fermionic component, along with a vector meson coupled to the Dirac spinor. The full expression is given by [71, 72, 73]
@@
where the strength tensor of the vector meson is $F_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$. The dark fermion mass is $m_{\chi}$, and $m_{V}$ is the mass of the dark vector meson. We do not include here the dark scalar meson, as done in [71], in order to simplify the model and avoid one more free parameter to be fixed. Besides that, the main effect for increasing the final compact star mass is due to repulsion interactions. Therefore, the inclusion of a scalar particle would add the attraction to the system and, consequently, a decreasing of the repulsion as a net feature.
In a similar way as done in relativistic hadronic models that present the same mathematical formulation for the Lagrangian density, we employ here the mean field approach in order to compute energy density and pressure of the dark sector of the system. These expressions read
@@
Despite efforts dedicated to the direct and indirect detection methods for dark matter in the last years [75], the true nature of DM particle is still not well established. This is the origin of the many possible candidates studied over the years, as the weakly interacting massive particles, gravitinos, axinos, axions, sterile neutrinos, WIMPzillas, supersymmetric Q-balls, and mirror matter [76, 77]. Among these options, both fermionic and bosonic dark particles are considered. In order to take into account the latter ones, we assume the bosonic asymmetric dark matter model proposed in [78] and explored in details in [79]. After neglecting interactions between the dark boson with visible matter, and approximating the spacetime to be flat, the Lagrangian density for the model reads
@@
$-\left.\frac{1}{4}Z_{\mu\nu}Z^{\mu\nu}\right)$ (17)
where $D_{\mu}=\partial_{\mu}+ig_{\sigma}\phi_{\mu}$, $g_{\sigma}$ is the interaction strength of the dark scalar complex field $\sigma$ with the dark vector field $\phi^{\mu}$, $Z_{\mu\nu}=\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}$, and $g$ is the determinant of the metric. The masses of the dark scalar and dark vector fields are $m_{\sigma}$ and $m_{\phi}$, respectively. Usual quantum field theory techniques, along with the mean field approximation, lead to the following energy-momentum tensor
@@
In the analysis of bosonic DM inside compact stars, it is important to emphasize that Black holes can be formed when the Chandrasekhar limit is achieved [83]. In the case of fermionic matter in ground state, the Pauli exclusion principle helps balancing gravity since degeneracy pressure is present. For bosonic systems, on the other hand, the lack of Fermi pressure sets this limit to a smaller value. However, it can be increased due to repulsive interactions like the one included in Eq. (17). In order to satisfy this condition, we restrict the values of $m_{\sigma}$ and $C_{\sigma\phi}$ to $10^{-2}~{}\mbox{MeV}\leqslant m_{\sigma}\leqslant 10^{8}~{}\mbox{MeV}$ and $10^{-2}~{}\mbox{MeV}^{-1}\leqslant C_{\sigma\phi}\leqslant 10^{3}~{}\mbox{MeV}^{-1}$, respectively, in agreement to the findings of [79].

### v3

@@ Abstract
In this work, we explore the possible scenario of strange stars admixed with fermionic or bosonic dark matter. For the description of the “visible” sector, we use a specific phenomenological quark model that takes into account in-medium effects for the quark masses through a suitable baryonic density dependence, in which the free parameters are chosen from the analysis of the stellar matter stability window (parametrizations presenting lower energy per baryon than iron nuclei). In the case of the dark sector, we investigate the predictions of fermionic and bosonic models. In the former, we consider a spin $1/2$ dark particle, and the latter is described by a dark scalar meson. Both models present a repulsive vector interaction, particularly important for the bosonic model since it helps avoid the collapse of the star due to the lack of degeneracy pressure. Our results point out possible stable strange stars configurations in agreement with data from the NASA NICER measurements of the pulsars PSR J0030+0451 and PSR J0740+6620, including the latest observational data, as well as the observed HESS J1731-347 mass-radius relationship. As another feature, we find stars with dark matter halo configurations for lower values of the dark particle mass.
@@ I Introduction
Stars are among the most studied astrophysical systems in the universe. However, their detailed composition, structure, and dynamics remain not fully understood by astrophysicists. For decades, these compact objects have been the subject of intense research, encompassing various areas of physics, such as thermodynamics, quantum field theory, general relativity, and nuclear reactions. The extreme conditions in the cores of stars, regions expected to be in a very high-density regime, also allow for the investigation of strongly interacting matter.
The theoretical input used to accurately describe compressed stellar matter (CSM) comes from models in which hadrons (baryons and mesons) are the fundamental degrees of freedom. Nevertheless, another branch of study considers quarks as the elementary constituents, using models that derive equations of state applied to the CSM formulation. This approach is based on the quark star hypothesis, proposed long ago in [1, 2], and later supported by the strange matter hypothesis [3, 4, 5, 6, 7, 8]: quark matter consisting of $u$, $d$, and $s$ quarks in beta equilibrium may have a lower energy per baryon than hadronic matter. Evidence for the presence of quark matter in the cores of massive stars has been recently discussed in Refs. [9, 10].
@@
Both hadronic and phenomenological quark models, developed for the aforementioned purpose, can be tested against a vast number of recent astrophysical observations. These new measurements have emerged primarily due to: (i) the detection of gravitational waves from a binary system, along with their respective electromagnetic counterparts [11], as observed by the LIGO and Virgo Collaboration [12, 13, 14]; (ii) data from massive millisecond pulsars PSR J0030+0451 [15, 16] and PSR J0740+6620 [17, 18, 19, 20, 21], provided by NASA’s Neutron Star Interior Composition Explorer (NICER), an X-ray telescope aboard the International Space Station [22]; and (iii) the first simultaneous measurements of mass, radius, and surface temperature of the compact star HESS J1731-347 [23].
@@
Recently, many researchers have dedicated efforts to better understanding the physics of these objects, which may also contain another mysterious component known as dark matter (DM) [24, 25, 26]. The main evidence for the existence of non-luminous matter comes from the analysis of galaxy rotation curves and gravitational lensing methods [27, 28], which indicate that the visible mass of galaxy clusters accounts for only about 10% to 20%, with the remaining mass attributed to DM. Measurements of the anisotropy of the cosmic microwave background also suggest the presence of DM [29, 30, 31, 32, 33, 34]. It is widely accepted that the universe, as we know it today, is composed of approximately 27% DM, 68% dark energy, and 5% ordinary matter.
@@
In this work, we investigate the possibility of the existence of strange stars admixed with fermionic and bosonic dark matter, in accordance with the astrophysical constraints mentioned earlier. To this end, we employ a specific effective quark model in the “visible” sector, namely, the equiparticle (EQP) model in which the constituent quark masses ($m_{u,d,s}$) are medium dependent, specifically density-dependent functions in the case of stellar matter at $T=0$. The effects of temperature on $m_{u,d,s}$ have been explored in [35, 36, 37, 38], while in [39], modified density dependencies of these quantities have been proposed. The thermodynamical inconsistency of the original EQP model was resolved in [40] by establishing a suitable connection between the Fermi momentum of the quark and the corresponding effective chemical potential. In [41], some of us further improved this model to incorporate deconfinement effects by introducing the traced Polyakov loop as the order parameter for the first-order confinement/deconfinement phase transition. This revised formulation was named the PEQM model.
@@
Here, we adopt the two-fluid approach, which considers only gravitational interaction between DM and quark matter. The calculations and results obtained in this investigation, all expressed in natural units ($\hbar=c=G=1$), are organized as follows. In Sec. II, we present the main quantities related to the effective EQP model and establish the conditions for the physical stability window. In Sec. III, we detail the main equations of state used to describe the dark sector in our analysis, including fermionic and bosonic DM models. The results concerning strange stars derived from this formalism are presented in Sec. IV. Finally, our concluding remarks are provided in Sec. V.
@@ II.1 Formalism
The EQP model was proposed in [40] as a baryonic density ($\rho_{b}$) and temperature ($T$) dependent model. However, here we focus on the zero temperature regime, as done in [42] and [41], for the study of cold quark stars. For this reason, the formalism of the model will be presented for this case. For an infinitely large system composed of up ($u$), down ($d$), and strange ($s$) quarks, the equations of state are obtained from fundamental thermodynamical relations. For the pressure we have
@@
where $m_{I}$ is the density-dependent quark mass term. The energy density reads
@@
Following up on this, for the EQP model the quark masses are constructed as being density-dependent as follows:
@@
where $m_{i0}$ is the current mass of the quark $i$. The interaction effects between the quarks are contained in the parameters $C$ and $D$, which are chosen from a stability window which will be described later. When working with compact objects, there is an interest in achieving higher stellar masses to satisfy the observational constraints. For this reason, the parameter $C$ plays an important role highlighting the necessity of incorporating this parameter into the mass scaling derived from [40] and presented in Eq. (10).
@@ II.2 Stellar matter
Given the interest in studying compact objects, it is necessary to establish the conditions for the matter present in its interior. At zero temperature and high-density regime, stellar matter in pure quark or hybrid stars [48, 49, 50, 51, 52, 53, 54, 55] can be described by effective quark models.
The most powerful mechanism of neutrino emission in stars is the Urca process [56, 57], responsible for their cooling since in this process stars lose a significant amount of thermal energy. The rate of neutrino emission influences how quickly the star cools down from its initial high temperature. After neutrinos are emitted from the star, its internal conditions undergo some changes, where the beta equilibrium takes place. This condition is given by $\mu_{u}+\mu_{e}=\mu_{d}=\mu_{s}$ in quark stars, which are the chemical potentials of the particles involved. In terms of the previously discussed model, this relation can be written as $\mu_{u}^{*}+\mu_{e}=\mu_{d}^{*}=\mu_{s}^{*}$, where the chemical potentials of the quarks are now the effective ones. The electron ($e$) does not have an effective chemical potential because its mass is constant. This equilibrium condition ensures that the star remains stable and maintains its composition over cosmic timescales. Furthermore, another condition has to be satisfied in compact stars, namely, the charge neutrality condition, given here by $\frac{2}{3}\rho_{u}-\frac{1}{3}\rho_{d}-\frac{1}{3}\rho_{s}-\rho_{e}=0$. These conditions exist due to the presence of leptons (electrons in this case), where weak interactions happen such as $d,s\leftrightarrow u+e+\bar{\nu}_{e}$. Because of the electrons’ presence, stellar matter is described by the sum of quarks and electrons’ contribution. Considering the (massless) electrons contribution, the total energy density and pressure of the part of the system composed of “visible” matter are
@@ II.3 Stability window
For the discussion of strange quark matter (SQM) in the context of quark stars, one has to consider the Bodmer-Witten hypothesis [3, 4]. It suggests that under certain conditions, such as at high densities or pressures, SQM may be the true ground state of matter rather than normal nuclear matter. Therefore, in order to establish the stability of SQM we need to investigate it against nuclear matter, which can be done by evaluating the minimum of the energy per baryon, $(E/A)_{\mbox{\tiny min}}$. This evaluation is then classified as follows: (i) SQM is stable when the minimum of the energy per baryon is lower than the binding energy of 56Fe (because it is the heaviest stable nucleus found in nature), i.e., $(E/A)_{\mbox{\tiny min}}\leqslant 930$ MeV; (ii) SQM is considered to be meta-stable when the minimum of the energy per baryon is higher than 930 MeV and lower than 939 MeV, which is the nucleon mass, i.e., 930 $930\mbox{ MeV}<(E/A)_{\mbox{\tiny min}}\leqslant 939$ MeV; (iii) SQM is unstable only if the minimum of the energy per baryon is higher than 939 MeV, i.e, $(E/A)_{\mbox{\tiny min}}>939$ MeV. In the case of stellar matter, we have $E/A=\mathcal{E}_{\mbox{\tiny vis}}/\rho_{b}$.
@@
In the figure, we observe different regions of stability for SQM. The pink one corresponds to the absolutely stable state of SQM, which is our area of interest. Below this area, we find the green region where two-flavor quark matter (2QM) is stable and where the SQM is forbidden to happen. In the orange region, the meta-stable state of SQM is expected to occur, and in blue we have the unstable SQM. It is important to mention that this analysis takes into consideration the data provided by PDG [58] regarding the ranges for the current quark masses, $m_{u0}=2.16^{+0.49}_{-0.26}$, $m_{d0}=4.67^{+0.48}_{-0.17}$ MeV and $m_{s0}=93.4^{+8.6}_{-3.4}$ MeV. For stellar matter the chosen values are $m_{u0}=2.16$ MeV , $m_{d0}=5.15$ MeV, and $m_{s0}=90$ MeV. For the pair $(C,D)$ we take $C=0.81$ and $D^{1/2}=127$ MeV, since it produces more massive stars when only quark matter is considered.
For the sake of completeness, we show in Fig. 2 the equation of state of the EQPM model, along with its respective sound speed, for the stellar (beta equilibrated) matter case.
Figure 2: Panel (a): dark matter pressure as a function of its respective energy density. Panel (b): dark matter sound speed. Fermionic and bosonic cases are represented by full and dashed lines, respectively.
Notice that Fig. 2b confirms that the model satisfies causality, i.e., it presents no superluminal states for the energy density range used in the work.
@@ II.4 Polyakov equiparticle model
The Polyakov equiparticle (PEQP) model was proposed in [59, 41] to study the confinement/deconfinement phase transition, possible to happen inside strange stars. In that article, some of us perform a few modifications in the original EQP model, which will be briefly described in this section. In this improved model, gluons and their corresponding strong interaction dynamics are phenomenologically described by the traced Polyakov loop, $\Phi$. When $\Phi=0$, the quarks are confined, and with $\Phi\rightarrow 1$ they achieve the deconfined phase. In the NJL model [60, 61, 62, 63, 64, 65], this quantity was introduced, at finite temperature [66], to generate the so-called Polyakov-Nambu-Jona-Lasinio (PNJL) model [67, 68, 69, 70, 71, 72, 73]. At first, the PNJL model loses its characteristics of describing the confinement/deconfinement phase transition at $T=0$, but in [74] the authors were able to solve this problem by taking the $b_{2}(T)$ function, present in the Polyakov potential, and modifying it to become a function that is also dependent on the quark density. Later on, the PNJL model suffered another improvement done in [75, 76, 77], where the motivation was to ensure that the coupling constants would vanish at the deconfinement phase, i.e. at $\Phi=1$.
All these observations concerning the improvement of the original NJL model were also implemented in the EQP model. More specifically, the parameters $C$ and $D$ were modified as follows, $C\rightarrow C^{\prime}(C,\Phi)=C(1-\Phi^{2})$, and $D\rightarrow D^{\prime}(D,\Phi)=D(1-\Phi^{2})$, which lead us to a new equation for the quark masses,
@@ III Dark matter models
With regard to the dark sector of our approach, we choose to investigate the capability of the DM model in satisfying the recent astrophysical constraints when admixed to the previously described quark matter, by implementing a fermionic and a bosonic model. Concerning the first one, we adopt, as a starting point, a Lagrangian density presenting a kinetic term (Dirac Lagrangian density) of a single fermionic component, along with a vector meson coupled to the Dirac spinor. The full expression is given by [78, 79, 80]
@@
where the strength tensor of the vector meson is $F_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$. The dark fermion mass is $m_{\chi}$, and $m_{V}$ is the mass of the dark vector meson. We do not include here the dark scalar meson, as done in [78], in order to simplify the model and avoid one more free parameter to be fixed. Besides that, the main effect of increasing the final compact star mass is due to repulsion interactions. Therefore, the inclusion of a scalar particle would add attraction to the system and, consequently, a decreasing of the repulsion as a net feature.
In a similar way as done in relativistic hadronic models that present the same mathematical formulation for the Lagrangian density, we employ here the mean-field approach in order to compute energy density and pressure of the dark sector of the system. These expressions read
@@
Despite efforts dedicated to the direct and indirect detection methods for dark matter in the last years [82], the true nature of DM particles is still not well established. This is the origin of the many possible candidates studied over the years, as the weakly interacting massive particles, gravitinos, axinos, axions, sterile neutrinos, WIMPzillas, supersymmetric Q-balls, and mirror matter [83, 84]. Among these options, both fermionic and bosonic dark particles are considered. In order to take into account the latter ones, we assume the bosonic asymmetric dark matter model proposed in [85] and explored in detail in [86]. After neglecting interactions between the dark boson with visible matter, and approximating the spacetime to be flat, the Lagrangian density for the model reads
@@
$-\left.\frac{1}{4}Z_{\mu\nu}Z^{\mu\nu}\right),$ (17)
where $D_{\mu}=\partial_{\mu}+ig_{\sigma}\phi_{\mu}$, $g_{\sigma}$ is the interaction strength of the dark scalar complex field $\sigma$ with the dark vector field $\phi^{\mu}$, $Z_{\mu\nu}=\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}$, and $g$ is the determinant of the metric. The masses of the dark scalar and dark vector fields are $m_{\sigma}$ and $m_{\phi}$, respectively. Usual quantum field theory techniques, along with the mean-field approximation, lead to the following energy-momentum tensor
@@
In the analysis of bosonic DM inside compact stars, it is important to emphasize that Black holes can be formed when the Chandrasekhar limit is achieved [90]. In the case of fermionic matter in the ground state, the Pauli exclusion principle helps balance gravity since degeneracy pressure is present. For bosonic systems, on the other hand, the lack of Fermi pressure sets this limit to a smaller value. However, it can be increased due to repulsive interactions like the one included in Eq. (17). In order to satisfy this condition, we restrict the values of $m_{\sigma}$ and $C_{\sigma\phi}$ to $10^{-2}~{}\mbox{MeV}\leqslant m_{\sigma}\leqslant 10^{8}~{}\mbox{MeV}$ and $10^{-2}~{}\mbox{MeV}^{-1}\leqslant C_{\sigma\phi}\leqslant 10^{3}~{}\mbox{MeV}^{-1}$, respectively, in agreement to the findings of [86].